OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2

OpenStudy (anonymous):

We know that : \[f(x) = 2x +1\] \[g(x) = 3x +1\] here no relation of the g(x) to question is given therefore it's of no use FInd \[2[f(4)]\] \[= 2[2 * 4 + 1] = 2[9] = 18\]

OpenStudy (anonymous):

MY ANSWERS ARE 81 34 AND 65 18 ISNT CORRECT

OpenStudy (anonymous):

X SQUARED

OpenStudy (anonymous):

is that \[x^2\]?

OpenStudy (anonymous):

YES I HAVE TO FIND THE [f(4)] 2.

OpenStudy (anonymous):

book's answer or you own worked out answer..itr can't return 3 outputs

OpenStudy (anonymous):

BOOK ANSWERS ADJAX

OpenStudy (anonymous):

what this [f(4)] 2??? squared???

OpenStudy (anonymous):

YES ARYAN

OpenStudy (anonymous):

if : \[f(x) = x^2 + 1\] you meant then it should be: \[2[f(4)] = 2[16 + 1] = 2[17] \]

OpenStudy (anonymous):

i thought that x2 = 2x ..please use latex or '^' symbol when denoting powers

OpenStudy (anonymous):

If f(x) = x^ 2 + 1 and g(x) = 3x + 1, find [f(4)] ^2. HERES THE QUESTION MY BAD